# Higgs-Confinement Transitions

### v2

@@ 6.2.2 Deforming to Zero-Density Higgs-Yukawa-QCD
Let us start with three-flavor QCD at very large $\mu_{B}\gg\Lambda_{\text{QCD}}$, so we are in the weakly-coupled phase reviewed in section 6.1 above.909090 We also assume that the quark masses $m_{u},m_{d},m_{s}\lesssim\Lambda_{\text{QCD}}$, as is the case in real QCD. Without this assumption some inequalities that appear in the subsequent argument have to be modified slightly. In particular, the composite Higgs field in (6.2) gets a CFL vev
@@
which acts on the index $i^{\prime}=1,2,3$ and is unrelated to the $SU(3)_{f}$ flavor symmetry acting on the fermions.919191 Before breaking $U(1)_{B}$, the symmetry group is $\left(U(3)_{f}\times U(3)_{f}^{\prime}\right)/\mathbb{Z}_{3}$, where the $\mathbb{Z}_{3}$ quotient is due to identification of the diagonal $\mathbb{Z}_{3}$ flavor symmetry with the center of the $SU(3)_{c}$ gauge group. If we fix unitary gauge and remove all the NGBs eaten by the $SU(3)_{c}$ gauge fields from ${\mathcal{H}}_{a}^{i}$, then all components of $h_{a}^{i}$ effectively become physical,
@@ 6.3.1 Unexpected Transitions in the QCD Phase Diagram
We have shown that SPT jumps can can enforce Higgs-confinement transitions in theories where they cannot be anticipated by symmetry-breaking considerations in the spirit of Landau. We have also shown that QCD at very large $U(1)_{B}$ chemical potential $\mu_{B}\gg\Lambda_{\text{QCD}}$ is a nontrivial SPT with $\theta_{g}=\theta_{f}=\pi$, as well as well as a $U(1)_{B}$ breaking superfluid. It is therefore natural to suspect that the SPT may trigger new, unexpected transitions in the QCD phase diagram as a function of $\mu_{B}$.929292 As before, we consider the theory at zero temperature, but in general new features on the $T=0$ boundary of the QCD phase diagram also have implications at finite temperature.
@@
As we now explain, we consider this picture to be in tension with the SPT phases at low and high densities that we have established: a trivial SPT in the confining phase at $\mu_{B}<\mu_{\text{SF}}$, and a non-trivial SPT with $\theta_{g}=\theta_{f}=\pi$ in the high-density CFL phase at $\mu_{B}\gg\Lambda_{\text{QCD}}$. The only way these SPTs are compatible with the quark-hadron continuity proposal of [19] is that if the $U(1)_{B}$ superfluid transition at $\mu_{B}=\mu_{\text{SF}}$ exactly coincides with the SPT jump from $\theta_{g},\theta_{f}=0$ to $\theta_{g},\theta_{f}=\pi$. While we cannot rule this out rigorously (after all, the transition occurs deep within the strong-coupling region), we find it implausible – essentially because the two phenomena are not intrinsically intertwined and their coincidence would seem like a finely-tuned accident.939393 It is well known that QCD with physical quark masses displays a number of finely tuned phenomena, e.g. the famously shallow binding of the deuteron. For this reason we cannot simply dismiss the apparent fine tuning mentioned above out of hand. To wit:
@@
The first inequality ensures that $\langle H\rangle=0$, so that $U(1)_{B}$ is unbroken and $M_{H}$ is the mass of the $H$-particle. The remaining inequalities follow from ${\mathsf{T}}$ and ${\mathsf{P}}$ symmetry,949494 Note the actions of ${\mathsf{C}}$, ${\mathsf{P}}$, and ${\mathsf{T}}$ in (3.12) on the baryons, ${\mathsf{C}}:({\mathcal{B}}_{\alpha})_{i}^{j}\leftrightarrow(\widetilde{\mathcal{B}}_{\alpha})^{i}_{j}~{},\qquad{\mathsf{P}}:({\mathcal{B}}_{\alpha})_{i}^{j}\to i(\widetilde{\mathcal{B}}^{\dagger{\dot{\alpha}}})_{i}^{j}~{},\qquad{\mathsf{T}}:({\mathcal{B}}_{\alpha})_{i}^{j}\to({\mathcal{B}}^{\alpha})_{i}^{j}~{}.$ (6.23) and the requirement that the $\mu_{B}=0$ confining phase is a trivial SPT.
For simplicity, we assume that $M_{H}<2M_{{\bf 8}\text{ baryon}}$, so that the $H$ cannot decay into baryon pairs.959595 The picture is only slightly modified if $M_{H}>2M_{{\bf 8}\text{ baryon}}$. In this case the $U(1)_{B}$-breaking transition occurs at $\mu_{B}\sim M_{{\bf 8}\text{ baryon}}<{1\over 2}M_{H}$ and is triggered by BCS pairing of the ${\bf 8}$ baryons in the channel with the quantum numbers of the $H$. In this case the $H$-particle has the smallest baryon number per unit mass and is expected to condense at $\mu_{B}=\mu_{\text{SF}}\sim{1\over 2}M_{H}$, triggering the $U(1)_{B}$-breaking transition. The fate of the SPTs depends on the strength of this first-order transition. If the transition is sufficiently strong, then the vev $\langle H\rangle$ jumps to sufficiently large values to ensure that the Majorana mass $\sim Y_{\bf 8}|\langle H\rangle|$ of the baryons can overwhelm their Dirac mass $M_{{\bf 8}~{}\text{baryon}}$. If this happens, half of the 16 Weyl fermions comprising the ${\bf 8}$ Dirac baryons experience a sign flip for their real mass eigenvalues. This does not affect the gravitational SPT, which remains
@@ 6.3.2 New Phenomena Inside Neutron Stars
Crudely speaking, neutron stars explore the QCD phase diagram as a function of $\mu_{B}$, starting with large $\mu_{B}$ in the high-density core of the star and dropping to smaller $\mu_{B}$ in the lower-density outer regions. Let us assume that there are neutron stars whose cores access the deconfined CFL Higgs phase at large $\mu_{B}$, which we now know to also support a $\theta_{g}=\pi$ SPT. As we march radially outward from the core, we will eventually encounter a phase transition at which the SPT jumps from $\theta_{g}=\pi$ in the interior to $\theta_{g}=0$ in the outer region (see figure 9). The two phases are separated by a transition layer (the green region in figure  9), whose width and surface tension depend on the details of the phase transition.969696 We continue to use a scheme in which the SPT in the outer confined regions (as well as outside the star) is trivial, but this is only for convenience. The SPT jump across the transition layer is scheme independent.

### v3

@@ 6.2.2 Deforming to Zero-Density Higgs-Yukawa-QCD
Let us start with three-flavor QCD at very large $\mu_{B}\gg\Lambda_{\text{QCD}}$, so we are in the weakly-coupled phase reviewed in section 6.1 above.939393 We also assume that the quark masses $m_{u},m_{d},m_{s}\lesssim\Lambda_{\text{QCD}}$, as is the case in real QCD. Without this assumption some inequalities that appear in the subsequent argument have to be modified slightly. In particular, the composite Higgs field in (6.2) gets a CFL vev
@@
which acts on the index $i^{\prime}=1,2,3$ and is unrelated to the $SU(3)_{f}$ flavor symmetry acting on the fermions.949494 Before breaking $U(1)_{B}$, the symmetry group is $\left(U(3)_{f}\times U(3)_{f}^{\prime}\right)/\mathbb{Z}_{3}$, where the $\mathbb{Z}_{3}$ quotient is due to identification of the diagonal $\mathbb{Z}_{3}$ flavor symmetry with the center of the $SU(3)_{c}$ gauge group. If we fix unitary gauge and remove all the NGBs eaten by the $SU(3)_{c}$ gauge fields from ${\mathcal{H}}_{a}^{i}$, then all components of $h_{a}^{i}$ effectively become physical,
@@ 6.3.1 Unexpected Transitions in the QCD Phase Diagram
We have shown that SPT jumps can can enforce Higgs-confinement transitions in theories where they cannot be anticipated by symmetry-breaking considerations in the spirit of Landau. We have also shown that QCD at very large $U(1)_{B}$ chemical potential $\mu_{B}\gg\Lambda_{\text{QCD}}$ is a nontrivial SPT with $\theta_{g}=\theta_{f}=\pi$, as well as well as a $U(1)_{B}$ breaking superfluid. It is therefore natural to suspect that the SPT may trigger new, unexpected transitions in the QCD phase diagram as a function of $\mu_{B}$.959595 As before, we consider the theory at zero temperature, but in general new features on the $T=0$ boundary of the QCD phase diagram also have implications at finite temperature.
@@
As we now explain, we consider this picture to be in tension with the SPT phases at low and high densities that we have established: a trivial SPT in the confining phase at $\mu_{B}<\mu_{\text{SF}}$, and a non-trivial SPT with $\theta_{g}=\theta_{f}=\pi$ in the high-density CFL phase at $\mu_{B}\gg\Lambda_{\text{QCD}}$. The only way these SPTs are compatible with the quark-hadron continuity proposal of [19] is that if the $U(1)_{B}$ superfluid transition at $\mu_{B}=\mu_{\text{SF}}$ exactly coincides with the SPT jump from $\theta_{g},\theta_{f}=0$ to $\theta_{g},\theta_{f}=\pi$. While we cannot rule this out rigorously (after all, the transition occurs deep within the strong-coupling region), we find it implausible – essentially because the two phenomena are not intrinsically intertwined and their coincidence would seem like a finely-tuned accident.969696 It is well known that QCD with physical quark masses displays a number of finely tuned phenomena, e.g. the famously shallow binding of the deuteron. For this reason we cannot simply dismiss the apparent fine tuning mentioned above out of hand. To wit:
@@
The first inequality ensures that $\langle H\rangle=0$, so that $U(1)_{B}$ is unbroken and $M_{H}$ is the mass of the $H$-particle. The remaining inequalities follow from ${\mathsf{T}}$ and ${\mathsf{P}}$ symmetry,979797 Note the actions of ${\mathsf{C}}$, ${\mathsf{P}}$, and ${\mathsf{T}}$ in (3.12) on the baryons, ${\mathsf{C}}:({\mathcal{B}}_{\alpha})_{i}^{j}\leftrightarrow(\widetilde{\mathcal{B}}_{\alpha})^{i}_{j}~{},\qquad{\mathsf{P}}:({\mathcal{B}}_{\alpha})_{i}^{j}\to i(\widetilde{\mathcal{B}}^{\dagger{\dot{\alpha}}})_{i}^{j}~{},\qquad{\mathsf{T}}:({\mathcal{B}}_{\alpha})_{i}^{j}\to({\mathcal{B}}^{\alpha})_{i}^{j}~{}.$ (6.23) and the requirement that the $\mu_{B}=0$ confining phase is a trivial SPT.
For simplicity, we assume that $M_{H}<2M_{{\bf 8}\text{ baryon}}$, so that the $H$ cannot decay into baryon pairs.989898 The picture is only slightly modified if $M_{H}>2M_{{\bf 8}\text{ baryon}}$. In this case the $U(1)_{B}$-breaking transition occurs at $\mu_{B}\sim M_{{\bf 8}\text{ baryon}}<{1\over 2}M_{H}$ and is triggered by BCS pairing of the ${\bf 8}$ baryons in the channel with the quantum numbers of the $H$. In this case the $H$-particle has the smallest baryon number per unit mass and is expected to condense at $\mu_{B}=\mu_{\text{SF}}\sim{1\over 2}M_{H}$, triggering the $U(1)_{B}$-breaking transition. The fate of the SPTs depends on the strength of this first-order transition. If the transition is sufficiently strong, then the vev $\langle H\rangle$ jumps to sufficiently large values to ensure that the Majorana mass $\sim Y_{\bf 8}|\langle H\rangle|$ of the baryons can overwhelm their Dirac mass $M_{{\bf 8}~{}\text{baryon}}$. If this happens, half of the 16 Weyl fermions comprising the ${\bf 8}$ Dirac baryons experience a sign flip for their real mass eigenvalues. This does not affect the gravitational SPT, which remains
@@ 6.3.2 New Phenomena Inside Neutron Stars
Crudely speaking, neutron stars explore the QCD phase diagram as a function of $\mu_{B}$, starting with large $\mu_{B}$ in the high-density core of the star and dropping to smaller $\mu_{B}$ in the lower-density outer regions. Let us assume that there are neutron stars whose cores access the deconfined CFL Higgs phase at large $\mu_{B}$, which we now know to also support a $\theta_{g}=\pi$ SPT. As we march radially outward from the core, we will eventually encounter a phase transition at which the SPT jumps from $\theta_{g}=\pi$ in the interior to $\theta_{g}=0$ in the outer region (see figure 9). The two phases are separated by a transition layer (the green region in figure  9), whose width and surface tension depend on the details of the phase transition.999999 We continue to use a scheme in which the SPT in the outer confined regions (as well as outside the star) is trivial, but this is only for convenience. The SPT jump across the transition layer is scheme independent.